INAF LBC+MODS queue observing and AO PISCES, 2012 June 18-19

Observer: C. Arcidiacono, R. Briguglio, K. Boutsia, M. Faccini, F. Vitali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC + PISCES

Summary

Clear skies at the beninng of the night but slightly windy (15m/s). At the second half of the night we had some gusts of 19m/s
At 7PM (LMT) we open the dome to take some Brg flats with Pisces.
The seeing was very variable, between 1.0" and 2.0" and it was out of spec for the LBC program. We used Pisces for the entire night.
At the beginning of the night we tried to do the Mesa program (HIP74596), but because of the variable seeing, the loop would not remain stable. No useful data for this project.
We then started the Fiore program (Mrk231) and we completed 30% of the requested time, before it reached 40 elevation that was considered too low to go on.
We spent the last part of the night doing the Testi program (BDp404126/V1686/V1318). Almost 80% of the project was completed before twilight.
Ks flats obtained during twilight.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Mesa Hip74596 FeII 5min started doing ADI but loop kept opening because of variable seeing - no useful data
Fiore Mrk231 Ks 10min + sky +calib 30% of requested time
Testi BDp404 J, H, Ks ~10min x filter + sky same project as V1686/V1318
Testi V1686 J, H, Ks 5min for J/H and 10min for Ks + sky same project as V1686/V1318
Testi V1318 Ks, H2, Br-g ~10min for Ks + 20min for Brg/H2 + sky 80% complete

Details

02:00 Brg flats with Pisces

03:21 send first preset - HIP74596 (Mesa)
  • failed because the "ccd47 is in wrong state"
  • we need to restart the CCD47

03:40 preset now passed with success and it communicates with the AOS that is trying to close the loop
  • the problem was with the WFS again - when we sent the first preset it got reconfigured ...

03:45 loop closed - start with science integration for HIP74596

03:57 loop open during an offset because of "skip frame"
  • send preset again - this time in ADI mode
  • active optics loop is not running - that was a warning
  • some confusion about the ADI mode...

04:10 start science observing in ADI mode with FeII
  • dimm seeing 1.2"
  • repeat autogain because the seeing is changing - loop is open

04:21 resend preset - failed again
  • dimm seeing 1.6"
  • the loop keeps opening - we give up on this target

04:31 we slew to a new target - Mrk231 (Fiore)
  • the seeing is really variable

04:40 sending an ACE preset - close loop
  • 04:50 start obtaining science frames

05:27 loop open but the preset is not cancelled
  • JC manually closes the loop and we go on with the image acquisition

06:20 slew to a calibration star for the Mrk231
  • obtaining images in open loop

06:40 slew to BDp404 - Testi

07:07 resend the preset because during an offset the loop opened

07:40 send preset to V1686 (still Testi - part of the mosaic)

07:55 loop opened
  • fixed - resend the preset

08:58 loop open

09:04 on target again - obtaining data
  • it took some time to center the source because the offset command did not work properly

09:34 start with science block

10:05 loop open - resend preset

11:09 the loop is open, but we let it continue since it is taking sky frames...

11:11 repeat preset to go on with the dither pattern

11:46 Sky flats in Ks

12:36 start darks (the telescope was moving for the first 7 frames of 15sec) - End of the night

-- KonstantinaBoutsia - 19 Jun 2012|
Topic revision: r7 - 22 Jun 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback